SHIKOKU SECTION BYLAWS

In all instances, the Institute of Electrical and Electronics Engineers (IEEE) Bylaws, Constitution, Policy and Procedures manual, and the Regional Activities Board (RAB)Bylaws and Operations Manual will prevail when there is a conflict between these documents and Section Bylaws.

ARTICLE I - Name and Territory

Section 1 This organization shall be known as the Shikoku Section of the IEEE.

Section 2 The territory of the Shikoku Section, as approved by RAB, includes the following: Ehime, Kagawa, Kochi, Tokushima prefectures, JAPAN.

ARTICLE II - Officers

Section 1 The elected officers of the Shikoku Section shall be

Chair  Vice Chair
Secretary  Treasurer

Section 2 The terms of office of the elected officers shall be for one (1) year.

Section 3 Terms of office will begin on 1 January but, in any case, the outgoing officers will continue until their successors are duly elected and take office. The consecutive of service shall not exceed two years.

Section 4 Any vacancy occurring during the years shall be filled by a majority vote of the Section Executive Committee.

Section 5 The duties of Section Officers shall be as described in IEEE Bylaws and in the IEEE Section Operations Guide.

ARTICLE III - Standing Committees

Section 1 The Standing Committees of the Section will be as follows:

Membership Development  Student Activities
Nominations  Professional Activities

Section 2 The Chair of the Standing Committees will be appointed by the Section Chair with the approval of the Section Executive Committee, and their terms will expire on the end of the calendar year.

Section 3 Each Committee Chair will appoint his/her Committee members, with the approval of the Section Executive Committee, and their terms will expire on the end of the calendar year.

ARTICLE IV - Management

Section 1 The management of the Section shall be by the Section Executive Committee which shall consist of the elected officers, the Past Section Chair and the following: Standing Committees Chairs.

Section 2 A majority of the Section Executive Committee shall constitute a quorum, provided that delegates elected by the Section membership are at least one greater in number than delegates appointed to the Committee.

Section 3 A majority of the Section Executive Committee present shall be necessary in the conduct of its business.

Section 4 Meetings of the Section Executive Committee ordinarily will be held monthly and will be called by the Section Chair or by a request of any three members of the Section Executive Committee.

Section 5 The fiscal year of the Section shall be the calendar year.

ARTICLE V - Nomination and Election of Officers

Section 1 A Nominating Committee consisting of four (4) members, not then officers of the Section, shall be appointed by the Section Chair with the approval of the Section Executive Committee.

Section 2 The nominations of the Nominating Committee will be announced to the Section membership and, following this, a minimum of 28 days allowed for additional nominations by petition. To be valid, the petition must be signed by 2 % or more voting members.

Section 3 If only one nomination is made for each office, the election will be made at the Annual Meeting. If additional nomination are made, election will be by ballot, mailed to the membership with the vote counted by a Tellers Committee appointed by the Section Chair.

Section 4 The timetable for this procedure is as follows:

Appointment of Nominating Committee:  by 31 July
Announcement of Nominations:  by 31 August
Close nominations by petition:  by 30 September
Hold election:  by 30 November
If required, a ballot shall be mailed prior to by 31 October

Section 5 A plurality of the votes cast shall be necessary for election.

ARTICLE VI - Business Meeting

Section 1 In order to transact business at a Section meeting, at least three (3) members must be present to constitute a quorum.

ARTICLE VII - Finances

Section 1 All expenditures of Section funds must be approved by Section Treasurer or Chair.

Section 2 Without prior authorization of the IEEE Executive Committee, Section funds can be used only for normal operations of the Section.

Section 3 The Treasurer shall be authorized to draw funds as approved by the Section Executive Committee.

ARTICLE VIII - Amendments

Section 1 Proposals for amendments to these Bylaws may originate in the Section Executive Committee or by a petition signed by twenty percent(20%) or more voting members.

Section 2 Amendments to or revocation of these Bylaws shall be in accordance with the IEEE Bylaws, RAB Bylaws and RAB Operations Manual.

Interim Shikoku Section Chair

Date: 29 January 1999